# TorchCP: A Python Library for Conformal Prediction

### v3

@@ A.3 Regression
NORABS (Vovk et al., 2005): The NORABS score $V(\bm{x},y):=\frac{|f(\bm{x})-y|}{\hat{\sigma}(\bm{x})}$ is a normalized variant of ABS. It uses models $f(\bm{x})$ and $\hat{\sigma}(\bm{x})$ trained to maximize the likelihood of the data with respect to $\mathbb{E}[Y_{\text{test}}|X_{\text{test}}=\bm{x}]$ and $\sqrt{\text{Var}[Y_{\text{test}}|X_{\text{test}}=\bm{x}]}$, respectively. Then the prediction set for a new instance $\bm{x}_{n+1}$ is constructed by $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x}_{n+1})-\widehat{Q}_{1-\alpha}\cdot\hat{\sigma}(\bm{x}_{n+1}),f(\bm{x}_{n+1})+\widehat{Q}_{1-\alpha}\cdot\hat{\sigma}(\bm{x}_{n+1})].$
CQR (Romano et al., 2019): The CQR score $V(\bm{x},y):=\max(f(\bm{x})_{\alpha/2}-y,\;y-$$f(\bm{x})_{1-\alpha/2})$ measures the maximum absolute deviation between the true value $y$ and the predicted quantile interval $[f(\bm{x})_{\alpha/2},f(\bm{x})_{1-\alpha/2}]$. Then the prediction set for a new instance $\bm{x}_{n+1}$ is constructed by: $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x})_{\alpha/2}-\widehat{Q}_{1-\alpha},\;f(\bm{x})_{1-\alpha/2}+\widehat{Q}_{1-\alpha}].$
CQRR (Sesia and Candès, 2020): The CQRR score extends the standard CQR method by introducing a scaling mechanism. This score normalizes the deviations between the true value $y$ and the predicted quantile interval $[f(\bm{x})_{\alpha/2},f(\bm{x})_{1-\alpha/2}]$ by the width of the interval itself. Formally, for a data pair $(\bm{x},y)$, the CQRR score is defined as $V(\bm{x},y):=\max\left(\frac{f(\bm{x})_{\alpha/2}-y}{\hat{s}_{\alpha}(\bm{x})},\;\frac{y-f(\bm{x})_{1-\alpha/2}}{\hat{s}_{\alpha}(\bm{x})}\right),$ where $\hat{s}_{\alpha}(\bm{x}):=f(\bm{x})_{1-\alpha/2}-f(\bm{x})_{\alpha/2}+\epsilon$ is the scaling factor, and $\epsilon>0$ is a small constant added to prevent division by zero. The prediction set is given by $\mathcal{C}(\bm{x}_{n+1})=\left[f(\bm{x}_{n+1})_{\alpha/2}-\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha}(\bm{x}_{n+1}),\;f(\bm{x}_{n+1})_{1-\alpha/2}+\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha}(\bm{x}_{n+1})\right].$
CQRM (Sesia and Candès, 2020): The CQRM score extends the CQRR score by introducing two scaling factors to independently normalize the deviations for the lower and upper bounds of the prediction interval. This method requires the model to output predictions at three quantile levels: $[\alpha/2,1/2,1-\alpha/2]$. The two scaling factors are defined as $\hat{s}_{\alpha,\text{lo}}(\bm{x}):=f(\bm{x})_{1/2}-f(\bm{x})_{\alpha/2}+\epsilon$ and $\hat{s}_{\alpha,\text{up}}(\bm{x}):=f(\bm{x})_{1-\alpha/2}-f(\bm{x})_{1/2}+\epsilon$. The non-conformity score for a data pair $(\bm{x},y)$ and prediction interval are then defined as:
@@
CQRFM (Kivaranovic et al., 2020): The CQRFM score is similar to CQRM, as it also requires the model to output predicted quantiles at $[\alpha/2,1/2,1-\alpha/2]$. However, it normalizes the deviations from the median $f(\bm{x})_{1/2}$ by the fractional distance to the nearest quantile bounds. The two fractional scaling factors are defined in the same way as in CQRM. The non-conformity score and prediction interval are defined as:
@@ A.4 GNN
We consider the prediction tasks on graph data. The Graph is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}:=\{v_{i}\}_{i=1}^{N}$ denotes the node set and $\mathcal{E}$ denotes the edge set with $|\mathcal{E}|=E$. Let $\bm{X}:=[\bm{x}_{1},\cdot\cdot\cdot,\bm{x}_{N}]^{T}$ be the node feature matrix, where $\bm{x}_{i}\in\mathbb{R}^{d}$ is a $d$-dimensional feature vector for node $v_{i}$. The label of $v_{i}$ is $y_{i}\in\mathcal{Y}$, where $\mathcal{Y}:=\{1,2,...,K\}$ denotes the label space.
DAPS (Zargarbashi et al., 2023): DAPS proposes a diffusion-based method that incorporates neighborhood information by leveraging the network homophily. Specifically, the score function is defined as: $V(\bm{x}_{i},y,\lambda):=(1-\lambda)\hat{V}(\bm{x}_{i},y)+\frac{\lambda}{|\mathcal{N}_{i}|}\sum_{v_{j}\in\mathcal{N}_{i}}\hat{V}(\bm{x}_{j},y),$ where $\hat{V}(\cdot,\cdot)$ is the basic non-conformity score function, e.g., APS, and $\lambda$ is a diffusion parameter.
@@ A.6 Metrics
In this section, we provide a brief overview of the evaluation metrics supported by TorchCP for various predictive tasks, as summarized in Table 4. These metrics facilitate rigorous and fine-grained assessment of prediction sets, offering deep insights into model uncertainty.
@@
Task Metrics
Classification Coverage, Size, CovGap (Ding et al., 2024), WSC (Romano et al., 2020), SSCV (Angelopoulos et al., 2020), Violated classes (Kasa and Taylor, 2023), DiffViolation (Angelopoulos et al., 2020), , etc.
Regression Coverage, Size
@@ C.1 Example codes for offline conformal prediction
    Classification. This example code [^7] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_splitcp_cifar100.py applies Split CP with the LAC score on CIFAR-100 using a pretrained ResNet-20. It evaluates the coverage and average set size of the resulting prediction sets.
@@
    Regression. The example code of regression [^8] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/regression_cqr_synthetic.py shows the complete workflow of applying Split CP with the Conformalized Quantile Regression score (Romano et al., 2019) on a synthetic dataset in regression task, including model training, calibration, prediction interval generation, and evaluation.
@@
    GNN. The node classification example [^9] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/gnn_transductive_coraml.py implements a Graph Convolutional Network (GCN) (Kipf and Welling, 2017) for node classification on the CoraML dataset (McCallum et al., 2000) under a transductive setting. Conformal prediction is then applied to generate prediction sets with guaranteed coverage.
@@
C.2 Example codes for online conformal prediction
Online conformal prediction means that after predicting a test point, its true label is immediately revealed, and the test sample is incorporated into the calibration set. TorchCP supports this setting as well, with the following examples:
@@ C.2 Example codes for online conformal prediction
    Regression. In TorchCP, the online setup of the regression task is integrated into ACIPredictor (Gibbs and Candes, 2021). Therefore, we added an example code  [^11] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/timeseries_aci_synthetic.py for the regression task in TorchCP.
@@
    Classification. For classification tasks, this online calibration approach can be naturally implemented in TorchCP by iteratively invoking the “calibrate()” function with new incoming data. To clarify this, we have added an online conformal prediction example 121212 https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_splitcp_cifar100_online.py for a classification task in TorchCP.
@@ D.2.2 Regression Task
Appendix E The Advantage of our modular design
The modular design of TorchCP prioritizes low coupling to ensure flexibility in integrating diverse training algorithms, non-conformity score functions, and prediction workflows, accommodating the wide variety of CP methods and tasks. In contrast, previous packages (e.g., PUNCC and Fortuna) focus on a monolithic design where predictors and scoring functions are often predefined within a specific predictor (e.g., Split CP). This limits extensibility when new CP predictors or custom score functions are needed. TorchCP ’s decoupled design allows users to independently select or define new trainers, scores, or predictors, enabling flexible integration of different components. In what follows, we elaborate on the concrete benefits of our modular design:
  • Flexible combination: users can freely combine Trainer, Score, and Predictor modules to customize conformal prediction workflows. For instance, a practitioner can use different non-conformity scores (e.g., LAC or APS) within any predictor or trainer, while other packages only implement scores for predictors.
  • Modification isolation: our module design ensures that changes to one module (e.g., refining the APS score) do not affect other modules. This isolation minimizes the risk of unintended side effects, reducing maintenance complexity and preventing errors from propagating across the system. For instance, tightly coupled systems like PUNCC require redefining the predictor to modify a scoring function, while TorchCP confines changes to the target module, streamlining updates and preserving stability.
  • Research extensibility: by adhering to standardized interfaces, new functionality (e.g., a novel non-conformity score) can be integrated into TorchCP without altering existing modules. This contrasts with PUNCC, where introducing new scoring methods demands extensive predictor modifications. Extensibility supports rapid iteration and long-term evolvability, enabling efficient adaptation to new research or applications.

### v4

@@ A.3 DNN Regressor
NORABS (Papadopoulos et al., 2011; Papadopoulos and Haralambous, 2011): The NORABS score $V(\bm{x},y):=\frac{|f(\bm{x})-y|}{\gamma+\xi(\bm{x})}$ normalizes the regression error using difficulty measure deviation $\xi(\bm{x})$. The prediction set for $\bm{x}_{n+1}$ is $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x}_{n+1})-\widehat{Q}_{1-\alpha}\cdot(\gamma+\xi(\bm{x}_{n+1})),f(\bm{x}{n+1})+\widehat{Q}_{1-\alpha}\cdot(\gamma+\xi(\bm{x}_{n+1}))]$. Additionally, we provide the normalized score for Gaussian Process Regression (Papadopoulos, 2024).
CQR (Romano et al., 2019): The CQR score $V(\bm{x},y):=\max(f(\bm{x})_{\alpha/2}-y,\;y$-$f(\bm{x})_{1-\alpha/2})$ measures the maximum absolute deviation between the true value $y$ and the predicted quantile interval $f(\bm{x})_{\alpha/2},f(\bm{x})_{1-\alpha/2}]$. Then the prediction set for a new instance $\bm{x}_{n+1}$ is constructed by: $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x})_{\alpha/2}-\widehat{Q}_{1-\alpha},\;f(\bm{x})_{1-\alpha/2}+\widehat{Q}_{1-\alpha}].$
CQRR (Sesia and Candès, 2020): The CQRR score extends the standard CQR method by introducing a scaling mechanism. This score normalizes the deviations between the true value $y$ and the predicted quantile interval $f(\bm{x})_{\alpha/2},f(\bm{x})_{1-\alpha/2}]$ by the width of the interval itself. Formally, for a data pair $(\bm{x},y)$, the CQRR score is defined as $V(\bm{x},y):=\max\left(\frac{f(\bm{x})_{\alpha/2}-y}{\hat{s}_{\alpha}(\bm{x})},\;\frac{y-f(\bm{x})_{1-\alpha/2}}{\hat{s}_{\alpha}(\bm{x})}\right),$ where $\hat{s}_{\alpha}(\bm{x}):=f(\bm{x})_{1-\alpha/2}-f(\bm{x})_{\alpha/2}+\epsilon$ is the scaling factor, and $\epsilon>0$ is a small constant added to prevent division by zero. The prediction set is given by $\mathcal{C}(\bm{x}_{n+1})=\left[f(\bm{x}_{n+1})_{\alpha/2}-\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha}(\bm{x}_{n+1}),\;f(\bm{x}_{n+1})_{1-\alpha/2}+\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha}(\bm{x}_{n+1})\right].$
CQRM (Sesia and Candès, 2020): The CQRM score extends the CQRR score by introducing two scaling factors to independently normalize the deviations for the lower and upper bounds of the prediction interval. This method requires the model to output predictions at three quantile levels: $\alpha/2,1/2,1-\alpha/2]$. The two scaling factors are defined as $\hat{s}_{\alpha,\text{lo}}(\bm{x}):=f(\bm{x})_{1/2}-f(\bm{x})_{\alpha/2}+\epsilon$ and $\hat{s}_{\alpha,\text{up}}(\bm{x}):=f(\bm{x})_{1-\alpha/2}-f(\bm{x})_{1/2}+\epsilon$. The non-conformity score for a data pair $(\bm{x},y)$ and prediction interval are then defined as:
@@
CQRFM (Kivaranovic et al., 2020): The CQRFM score is similar to CQRM, as it also requires the model to output predicted quantiles at $\alpha/2,1/2,1-\alpha/2]$. However, it normalizes the deviations from the median $f(\bm{x})_{1/2}$ by the fractional distance to the nearest quantile bounds. The two fractional scaling factors are defined in the same way as in CQRM. The non-conformity score and prediction interval are defined as:
@@ A.4 GNN
In this section, we discuss the applications of conformal prediction algorithms on graph data. The Graph is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}:=\{v_{i}\}_{i=1}^{N}$ denotes the node set and $\mathcal{E}$ denotes the edge set with $|\mathcal{E}|=E$. Let $\bm{X}:=[\bm{x}_{1},\cdot\cdot\cdot,\bm{x}_{N}]^{T}$ be the node feature matrix, where $\bm{x}_{i}\in\mathbb{R}^{d}$ is a $d$-dimensional feature vector for node $v_{i}$. The label of $v_{i}$ is $y_{i}\in\mathcal{Y}$, where $\mathcal{Y}:=\{1,2,...,K\}$ denotes the label space.
[…]
@@ A.6 Metrics
In this section, we provide a brief overview of the evaluation metrics supported by TorchCP for diverse modern models, as summarized in Table 4. These metrics facilitate rigorous and fine-grained assessment of prediction sets, offering deep insights into model uncertainty.
@@
Deep Models Metrics
DNN Classifier Coverage, Size, CovGap (Ding et al., 2024), WSC (Romano et al., 2020), SSCV (Angelopoulos et al., 2020), Violated classes (Kasa and Taylor, 2023), DiffViolation (Angelopoulos et al., 2020), S Criterion, N Criterion, U Criterion, F Criterion, M Criterion, F Criterion, OU Criterion, OF Criterion, OM Criterion, OE Criterion (Vovk et al., 2022), etc.
DNN Regressor Coverage, Size
@@ C.1 Example codes for offline conformal prediction
    DNN classifier. This example code [^10] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_splitcp_cifar100.py applies Split CP with the LAC score on CIFAR-100 using a pretrained ResNet-20. It evaluates the coverage and average set size of the resulting prediction sets.
@@
    DNN Regressor. The example code of regression [^11] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/regression_cqr_synthetic.py shows the complete workflow of applying Split CP with the Conformalized Quantile Regression score (Romano et al., 2019) on a synthetic dataset in a regression task, including model training, calibration, prediction interval generation, and evaluation.
@@
    GNN. The GNN example [^12] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/gnn_transductive_coraml.py implements a Graph Convolutional Network (GCN) (Kipf and Welling, 2017) for node classification on the CoraML dataset (McCallum et al., 2000) under a transductive setting. Conformal prediction is then applied to generate prediction sets with guaranteed coverage.
@@
C.2 Example codes for Semi-online conformal prediction
Semi-online conformal prediction means that after predicting a test point, its true label is immediately revealed, and the test sample is incorporated into the calibration set. TorchCP supports this setting as well, with the following examples:
@@ C.2 Example codes for Semi-online conformal prediction
    DNN regressor. In TorchCP, the semi-online setup of the regression task is integrated into ACIPredictor (Gibbs and Candes, 2021). Therefore, we added an example code  [^14] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/timeseries_aci_synthetic.py for the regression task in TorchCP.
@@
    DNN Classifier. For DNN classification tasks, this semi-online calibration approach can be naturally implemented in TorchCP by iteratively invoking the “calibrate()” function with new incoming data. To clarify this, we have added a semi-online conformal prediction example 151515 https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_splitcp_cifar100_online.py for a classification task in TorchCP.
@@ D.2.1 Classification Task
D.2.2 Regression Task
Appendix E The Advantage of our low-coupling design
Low-coupling design is one of the most fundamental concepts in object-oriented design, which minimizes the dependency a class has on other classes. The low-coupling design of TorchCP ensures flexibility in integrating diverse training algorithms, non-conformity score functions, and prediction workflows, accommodating the wide variety of CP methods and tasks. In contrast, most current packages focus on a high-coupling design where predictors and scoring functions are often predefined within a specific predictor (e.g., Split CP). This limits extensibility when new CP predictors or custom score functions are needed. TorchCP ’s decoupled design allows users to independently select or define new trainers, scores, or predictors, enabling flexible integration of different components. In what follows, we elaborate on the concrete benefits of our low-coupling design:
  • Flexible combination: users can freely combine Trainer, Score, and Predictor modules to customize conformal prediction workflows. For instance, a practitioner can use different non-conformity scores (e.g., LAC or APS) within any predictor or trainer, while other packages only implement scores for predictors. Moreover, some libraries (e.g., TorchUQ) implemented various non-conformity scores as functions, but this style suffers from extending to those scores with extra hyperparameters (e.g., the $k$ and $\lambda$ of RAPS).
  • Modification isolation: our low-coupling design ensures that changes to one module (e.g., refining the APS score) do not affect other modules. This isolation minimizes the risk of unintended side effects, reducing maintenance complexity and preventing errors from propagating across the system. For instance, tightly coupled systems like PUNCC require redefining the predictor to modify a scoring function, while TorchCP confines changes to the target module, streamlining updates and preserving stability.
  • Research extensibility: by adhering to standardized interfaces, new functionality (e.g., a novel non-conformity score) can be integrated into TorchCP without altering existing modules. Differently, MAPIE follows a high-coupling design, making it challenging to introduce new predictors, like class-conditional CP. Extensibility supports rapid iteration and long-term evolvability, enabling efficient adaptation to new research or applications.
@@ Appendix E The Advantage of our low-coupling design
Appendix F Comparison with related libraries
Numerous libraries implement conformal prediction (CP) algorithms, yet their suitability for deep learning (DL) scenarios differs significantly, particularly in supporting modern DL models like GNNs and LLMs, as well as their scalability for large-scale datasets. In Table 8, we present a concise comparison of open-source CP libraries, focusing on supported model types and functionalities:
  • Model type refers to the types of machine learning models the library can work with. They are classified based on differences in input and output data, which influence the selection of applicable CP algorithms.
  • Functionality describes the operational capabilities of the library, such as how it processes data, computational optimization, or prediction forms.
In the following, we compare several popular CP libraries, emphasizing TorchCP’s superior compatibility and performance with DL applications.
nonconformist (Linusson et al., 2020): A NumPy-based library supporting basic CP algorithms for classification and regression. It lacks GPU acceleration, batch processing, and compatibility with DL models like DNN Classifier, DNN Regressor, GNNs, or LLMs, making it unsuitable for complex DL scenarios. While nonconformist implements predictors and scores as different classes following a low-coupling design, its algorithmic coverage is limited, implementing only inductive CP and aggregated CP with 5 scores.
TorchUQ (Zhao et al., 2024): A PyTorch library focused on uncertainty quantification, including some CP methods for regression, with GPU support. However, it lacks comprehensive algorithm implementation, omitting CP algorithms for classifiers and CP-specific training, and does not support advanced deep learning models like GNNs or LLMs, limiting its applicability for complex deep learning Scenarios. TorchUQ uses a dictionary to map names to score functions in predictors, enabling developers to define new scores without modifying predictor logic. However, this style struggles to accommodate scores with additional hyperparameters (e.g., $k$ and $\lambda$ in RAPS), requiring extra logic to manage these variations, which increases coupling and complexity.
Crepes (Boström, 2022): a NumPy-based library with extensive prediction forms that supports conformal classifiers, regressors, and predictive systems with semi-online CP, excelling in statistical classification and regression tasks. However, it lacks GPU acceleration, batch processing, and seamless extensions to modern models like GNNs or LLMs, limiting scalability for engineers tackling large-scale or complex deep learning problems. Similar to TorchUQ, Crepes implements non-conformity scores as functions passed to predictors, so it also suffers from supporting scores with additional hyperparameters (e.g., $k$ and $\lambda$ in RAPS).
PUNCC (Mendil et al., 2023): A scikit-learn-compatible library delivers Conformal Prediction for tasks like regression, classification, and anomaly detection, but struggles with modern DL applications. It does not support advanced DL models such as DNN, GNNs, or LLMs, lacks GPU acceleration and batch processing, and performs significantly slower on large datasets (up to 10x slower than TorchCP in ImageNet, see Figure 6(b)). PUNCC defines nonconformity scores as functions, but ties each score to a dedicated predictor, resulting in high coupling.
MAPIE (Cordier et al., 2023): A scikit-learn-based library providing CP sets/intervals for classical classification/regression tasks. It excels in traditional machine learning workflows but lacks support for deep learning models, such as DNNs, GNNs, or LLMs. It also does not offer online CP, GPU acceleration, and batch processing. MAPIE implements scores as different classes with a predict function, making it challenging to support other predictors (like class-conditional CP (Vovk, 2012)).
Fortuna (Detommaso et al., 2024): a JAX-based library, supports CP methods in deep learning with GPU acceleration and select conformal prediction methods. It enhances scalability for deep learning but has limited algorithmic coverage, lacking training algorithms and online CP, and does not support GNNs or LLMs. Additionally, JAX’s steeper learning curve may challenge statisticians familiar with PyTorch. Like PUNCC, Fortuna defines a dedicated classifier for each score, leading to high-coupling design.
In contrast, TorchCP, a PyTorch-native library, significantly enhances conformal prediction for deep learning by providing a comprehensive suite of state-of-the-art CP algorithms. It offers exceptional extensibility through low-coupling design with three modules (See Appendix E) and unified interfaces across various algorithms (See Appendix B). Notably, TorchCP seamlessly integrates CP algorithms with advanced deep learning models, such as GNNs and LLMs, and leverages GPU-accelerated batch processing to ensure scalability for large-scale deep learning scenarios (See Appendix D.2).